# Flat Spectrum Radio Quasars as high-energy neutrino sources

### v1

@@
Flat Spectrum Radio Quasars as high-energy neutrino sources
The astrophysical sources responsible for the production of high energy neutrinos remain largely uncertain. The strongest associations suggest a correlation between neutrinos and active galactic nuclei (AGN). However, it is still unclear which specific regions and mechanisms of the accreting supermassive black hole (SMBH) are responsible for their production. In this paper we investigate the correlation between the positions of IceCat-1 neutrino events and a large, optically selected QSO catalog extracted from the Sloan Digital Sky Survey. Within this sample, we distinguish radio-quiet QSOs from flat-spectrum radio QSOs (FSRQs) based on radio emission data from the Cosmic Lens All Sky Survey (CLASS) catalog. While the associations between neutrino events and radio-quiet QSOs are consistent with being all random matches, FSRQs exhibit a moderately significant correlation (2.7) with neutrino positions. Additionally we observe that the distribution of minimum distances between neutrino events and FSRQs differs significantly for events at declinations above and below 20∘. In particular, using the KS test, we find that the high-declination event distribution shows a strong deviation (4) from a random distribution. We interpret all these results as an indication that a large fraction (60%) of the neutrino events observed by IceCube could be produced by the FSRQ and that the emission mechanism is likely related to the relativistic jets rather than the ”radio-quiet” (RQ) component of these sources, such as the accretion disk or corona.
@@
Active Galactic Nuclei (AGN) are commonly considered as promising candidates for the emission of the high energy astrophysical neutrinos that are routinely detected by the IceCube experiment since 2013 (IceCube Collaboration 2013 for a review). Indeed, AGN can be powerful accelerators of protons, either through relativistic jets or through the accretion disk/hot corona, and these particles can then interact with the intense radiation fields present around the active nucleus producing cascades of particles including energetic neutrinos. The emission from relativistic jets pointing towards Earth, in particular, can be very strong due to relativistic boosting (beaming). This is the reason why these ”oriented” sources, called blazars, have been extensively investigated in the last 10 years to assess their possible relevance for the neutrino astrophysics (Dermer et al. 2014b; Tavecchio et al. 2014; Tavecchio & Ghisellini 2015; Giommi et al. 2020a, b; Giommi & Padovani 2021; Plavin et al. 2020, 2023; Buson et al. 2019, 2022, 2023; Bellenghi et al. 2023).
@@
Blazars can be observed in a wide variety of “flavours” depending on the shape of the Spectral Energy Distribution (SED) and/or to the optical properties. In particular, blazars in which the synchrotron emission peaks at low (1014 Hz), intermediate (1014-1015 Hz) or high (1015 Hz) frequencies are called low-intermediate-high frequency peak sources (LSP, ISP, HSP) respectively (Padovani & Giommi 1995; Abdo et al. 2010). Based on their optical properties, instead, blazars with featureless spectra are dubbed as BL Lacertae (BL Lac) objects while they are called Flat spectrum Radio Quasars (FSRQ) when strong and broad emission lines, similar to those usually found in non-jetted AGN, are observed. The distinction between BL Lac and FSRQ could be relevant for the neutrino physics, since the very existence of a Broad Line Region (BLR) and of a “standard” accretion disk is directly connected to the presence of the radiation field necessary for an efficient neutrino emission (e.g. Dermer et al. 2014a). For this reason, FSRQ have been originally considered as the most promising neutrino emitters, although also BL Lacs could be viable candidates if velocity structures are present in the jets (e.g. Tavecchio et al. 2014; Tavecchio & Ghisellini 2015).
@@
Interestingly, the first firm (3) association between a neutrino event and an extragalactic source, based on the detection of flares temporarily coincident with neutrino arrival, and then confirmed by an excess of events in the archival IceCube data, was the source TXS 0506+056 (IceCube Collaboration et al. 2018b, a; Aartsen et al. 2020), an ISP which was originally classified as BL Lac object. However, a detailed analysis of the optical spectrum of this source has revealed the presence of a “standard” accretion disk emission, as in FSRQ, whose light was probably swamped by the strong non-thermal continuum coming from the oriented relativistic jet (Padovani et al. 2019). This observation puts TXS 0506+056 apart from “true” BL Lac objects that genuinely lack of an efficient accretion disk and of a BLR, probably due to their low accretion rate of the central Super Massive Black-hole (e.g. see Hardcastle & Croston 2020), and suggests an emission mechanism that is similar to that considered for FSRQ.
@@
The search for positional coincidences between neutrino events (or excesses of events) and extragalactic sources is hampered by the large positional uncertainties of the events that make the number of chance coincidences very high, even for relatively rare objects like blazars.
@@
Hints of positional correlations between neutrino events and BL Lac objects have been found by Padovani et al. (2016) and Giommi et al. (2020b). Other works suggest a possible connection of IceCube events with the class of blazars, including both FSRQ and BL Lacs (Buson et al. 2022, 2023), or with the radio brightest (150 mJy) blazars, again, including both classes (Plavin et al. 2020, 2021, 2023).
@@
Even if relativistic jets are natural candidates for neutrino emission, the recent association of a radio-quiet AGN (NGC1068) with neutrino events (IceCube Collaboration et al. 2022) suggests that other possible mechanisms may be at work. For instance, Abbasi et al. (2022) have shown that the acceleration of cosmic-rays responsible for the neutrino emission may occur in the accretion disk region of AGN. The multi-wavelength analysis of NGC 1068 conducted by Padovani et al. (2024), within a p- scenario, indicates that the most plausible location for the neutrino source is in close proximity to the SMBH, embedded in an extremely photon-dense environment. This finding suggests the AGN corona as the likely origin, while star-forming regions, the jet and the outflow are disfavored. Yang et al. (2025) find that even in the case of TXS0506+056, the accretion flow might explain both the long-term neutrino steady emission and the outburst, for which a super Eddington accretion is required. For this reason, the simple positional match of a sample of blazars with neutrino events, even when statistically significant, does not necessarily imply that the emission is due to the relativistic jets, since some of these sources (the FSRQ) also have a powerful accretion disk similar to that observed in radio-quiet QSO.
@@
In this paper we test the hypothesis that FSRQ could be responsible, at least in part, of the astrophysical neutrino events detected by IceCube and, at the same time, we aim at establishing whether this is really due to the presence of an oriented jet or, instead, can be associated to the radio-quiet component of a FSRQ i.e. the accretion disk and corona. To this end, we start from an optically selected sample of AGN (including both radio-loud and radio-quiet objects) and from this we extract both the FSRQ and the radio-quiet QSO samples. From the analysis of their correlation with neutrino events we expect to assess not only the association of neutrino production with QSOs, but also a direct probe on the possible origin site of the neutrinos (jets or disk-corona accretion ).
@@
In Section 2, we present the QSO data sample and the list of IceCube events that we consider for the analysis. In Section 3 we describe the positional QSOs-neutrinos cross-match and present the results for both FSRQ and RQ. In Section 4 we analyze the distribution of the minimum distance between neutrino and QSOs. In Section 5 the most-likely neutrino emitter candidates are presented individually while, in Section 6, we discuss the purity of the sample together with the estimated fraction of high energy neutrino events produced by FSRQ. Finally in Section 7, we put the results in the context of the recent findings presented in the literature.
@@
(*) Note that in the Ice-Cat-1 catalog two events named IC120523A are present and the one listed here is different from the one associated to 3C454.3 by Abbasi et al. (2023a).
@@
Throughout the paper, we assume a flat CDM cosmology with H0=71 km s-1 Mpc-1, =0.7 and =0.3. Spectral indices are given assuming S.
2 Quasar and neutrino catalogs
2.1 The QSO samples
@@
We start with the sample of type 1 AGN presented in Shen et al. (2011) (S11), which is based on the data from the Sloan Digital Sky Survey (SDSS). The sample includes 105,783 spectroscopically confirmed type 1 AGN, mostly in the northern hemisphere, and contains several pieces of information on the central engine, such as the mass of the SMBH and its accretion rate, which can be used to further investigate potential neutrino emitters. For our analysis we only consider the sky area within, 120 R.A. 260∘ and 0∘ Dec 60∘, where the coverage is uniform (Fig. 1) and that includes a total of 89,437 QSOs.
@@
As far as radio information is concerned, we use the Cosmic Lens All Sky Survey (CLASS, Myers et al. 2003; Browne et al. 2003) which homogeneously covers all the extragalactic Northern sky. The CLASS catalogue contains 11,000 flat spectrum radio sources selected at 5 GHz down to a well-defined flux density limit (30 mJy). Since blazars are flat spectrum radio sources, this survey represents the best starting point for our selection. The survey is also deep enough to select most of the blazars with an optical magnitude brighter than 20 (which roughly corresponds to the limit of the S11 sample). Thanks to the excellent positional accuracy of the survey (1″) the optical/radio association can be performed with high reliability (Fig. 1). The cross-correlation between the S11 sample and CLASS, produces 1218 sources (1122 in the region of interest), using 1″ positional tolerance. We call this sample CLASH (CLASS-SHEN) and throughout the rest of the paper, unless otherwise specified, flux densities are measured at 5 GHz.
@@
For comparison, we also build a sample of radio-quiet (RQ) QSOs, defined as those sources from S11 not detected in CLASS and located more than 5 away from any source in the FIRST catalog. FIRST is a radio survey at 1.4 GHz with flux density limit of 1 mJy, significantly deeper than CLASS and with a good positional accuracy ( 2). Given the typical magnitudes of the Shen et al. (2011) QSO, the non-detection of a source in the FIRST survey implies a radio-loudness111The radio-loudness is defined as the ratio between the radio and the optical monochromatic luminosities (per unit of frequency) computed, respectively, at 5 GHz and at 4400Å (Kellermann et al. 1989) below 10 i.e. in the typical radio-quiet regime (Kellermann et al. 1989). Therefore, the “radio-quiet” QSO sample should mostly contain non-jetted objects. This control sample comprises 73,339 “radio-quiet” objects.
@@
As shown in Fig. 2, the two populations are indistinguishable in terms of redshift, mass, bolometric luminosity and Eddington ratio. The only difference is the presence of a strong radio emission which we assume is produced by a relativistic jet oriented toward our line of sight.
@@ 2.2 The IceCube neutrino events
For our study we use the IceCube Event Catalog of Alert Tracks (IceCat-1, Abbasi et al. 2023b) which includes 267 non-vetoed track-like neutrino events likely produced by astrophysical sources between 2011 and 2021. Since the positional uncertainty is extremely large in a significant fraction of the events, we consider only those events with the best positional accuracy:
@@
We stress that these requirements were frozen at the beginning of the analysis, and no optimization has been performed. In the region of interest defined above (Fig. 1), these constraints produce a list of 15 events (Tab. 1), with IC121103A, IC190503 and IC200117A excluded because too close to the border. Following the approach of Plavin et al. (2023), also discussed in Abbasi et al. (2023a), the 90% error region is assumed to be encircled by the 4 elliptical segments with the RA and DEC errors as semi-axes. The energy distribution of the 15 events is shown in Fig. 3: the selection on the positional accuracy includes events belonging to the bulk of the population detected in the Northern sky with energies in the 100-500 Tev interval (Tab.1).
3 QSO / neutrino associations
3.1 CLASH-FSRQ / neutrino
@@
Considering the list of 15 IceCat-1 events and the 1122 CLASH FSRQ we find that 9 CLASH sources fall within the error boxes of 7 IceCUBE events (Tab. 1). Varying the flux density limit of the CLASH sample, the number of objects varies from 1122 with a limit of 30 mJy to 270, when the flux density limit is 200 mJy. The number of neutrino events, with at least one CLASH QSO within the 90% error region, varies from 7 to 3, while the number of CLASH included in the neutrino 90% conf. error region varies from 9 to 3 (Tab. 2).
@@
Multiple matches clearly indicate that the probability of chance coincidence is not negligible. We estimate the expected number of chance coincidences by randomly shifting QSO positions within a circular annulus of radius 3–10∘ around their original locations. This procedure was repeated 105 times for different radio flux density limits of the FSRQ sample. The statistical significance of the observed correlation (p-value) was calculated by counting the realizations in which the number of neutrino including at least a CLASH FSRQ exceeded that observed in the actual data set. The output of this Montecarlo (MC) simulation at different radio flux density limits are reported in Fig. 4 and in Tab. 2.
@@
These values show a minimum located at a flux density limit of 100 mJy (Fig. 4). At this flux density limit, the number of CLASH FSRQs is 497, with 7 of them in the error box of 6 neutrino events. The average number of associations of an equal number of random positions with the 15 neutrino error box is 1.3. We find that in 52 cases out of 1e5 an equal or larger number neutrino have a CLASH in the error box. Considering the number of different trials involved in this analysis (6 with different flux density limits), we estimate that the final probability corresponds to a significance of 2.7. We obtain identical results when repeating the simulation by randomly shifting the positions of the neutrinos instead of the QSOs.
fmin | QSOs | #n | #n | p-val | KS | KS (20∘) |
30 | 1122 | 7 | 2.6 | 6.3e-03 | 2.8e-02 | 8.2e-05 |
75 | 614 | 6 | 1.4 | 1.7e-03 | 6.5e-03 | 7.8e-06 |
100 | 497 | 6 | 1.2 | 5.3e-04 | 3.3e-03 | 4.8e-06 |
125 | 402 | 5 | 0.9 | 1.5e-03 | 2.4e-03 | 9.4e-06 |
175 | 300 | 3 | 0.7 | 3.0e-02 | 8.4e-03 | 5.3e-06 |
200 | 270 | 3 | 0.6 | 2.3e-02 | 2.9e-02 | 1.4e-04 |
3.2 Radio-quiet / neutrino

### v2

# Flat-spectrum radio quasars as high-energy neutrino sources

@@
Flat-spectrum radio quasars as high-energy neutrino sources
The astrophysical sources responsible for the production of high-energy neutrinos remain largely uncertain. The strongest associations suggest a correlation between neutrinos and active galactic nuclei. However, it is still unclear which specific regions and mechanisms of the accreting supermassive black hole are responsible for their production. In this paper we investigate the correlation between the positions of IceCat-1 neutrino events and a large, optically selected quasar catalogue extracted from the Sloan Digital Sky Survey. Within this sample, we distinguish radio-quiet quasars from flat-spectrum radio quasars (FSRQs) based on radio emission data from the Cosmic Lens All Sky Survey (CLASS) catalogue. While all the associations between neutrino events and radio-quiet quasars are consistent with being random matches, FSRQs exhibit a moderately significant correlation (2.7) with neutrino positions. Additionally, we observe that the distribution of minimum distances between neutrino events and FSRQs differs significantly for events at declinations above and below 20∘. In particular, using the Kolmogorov-Smirnov test, we find that the high-declination event distribution deviates strongly (4) from a random distribution. We interpret all these results as an indication that a large fraction of the neutrino events (60%) observed by IceCube could be produced by the FSRQs and that the emission mechanism is likely related to the relativistic jets rather than the radio-quiet component of these sources, such as the accretion disk or corona.
@@
Active galactic nuclei (AGNs) are commonly considered promising candidates for the emission of the high-energy astrophysical neutrinos that have routinely been detected by the IceCube experiment since 2013 (IceCube Collaboration 2013 for a review). Indeed, AGNs can be powerful accelerators of protons, either through relativistic jets or through the accretion disk, hot corona. These particles can then interact with the intense radiation fields present around the active nucleus, thereby producing cascades of particles, including energetic neutrinos. The emission from relativistic jets pointing towards Earth in particular can be very strong due to relativistic boosting (beaming). This is the reason why these ‘oriented’ sources, called blazars, have been extensively investigated in the last 10 years, to assess their possible relevance for neutrino astrophysics (Dermer et al. 2014b; Tavecchio et al. 2014; Tavecchio & Ghisellini 2015; Giommi et al. 2020a, b; Giommi & Padovani 2021; Plavin et al. 2020, 2023; Buson et al. 2019, 2022, 2023; Bellenghi et al. 2023).
@@
Blazars can be observed in a wide variety of ‘flavours’ depending on the shape of the spectral energy distribution and/or the optical properties. In particular, blazars in which the synchrotron emission peaks at low (1014 Hz), intermediate (1014-1015 Hz), or high (1015 Hz) frequencies are called low-, intermediate-, and high-frequency peak sources, respectively (Padovani & Giommi 1995; Abdo et al. 2010). Based on their optical properties, blazars with featureless spectra are instead dubbed BL Lacertae (BL Lac) objects; they are called flat spectrum radio quasars (FSRQs) when strong and broad emission lines, similar to those usually found in non-jetted AGNs, are observed. The distinction between BL Lac and FSRQs could be relevant for neutrino physics since the very existence of a broad-line region (BLR) and a ‘standard’ accretion disk is directly connected to the presence of the radiation field necessary for an efficient neutrino emission (e.g. Dermer et al. 2014a; Murase et al. 2014). For this reason, FSRQs were originally considered the most promising neutrino emitters, although BL Lacs could also be viable candidates if velocity structures are present in the jets (e.g. Tavecchio et al. 2014; Tavecchio & Ghisellini 2015).
@@
Interestingly, the first firm association (3) between a neutrino event and an extragalactic source, based on the detection of flares temporarily coincident with a neutrino’s arrival and then confirmed by an excess of events in the archival IceCube data, was the source TXS 0506+056 (IceCube Collaboration et al. 2018b, a; Aartsen et al. 2020), an intermediate-frequency peak source that was originally classified as a BL Lac object. However, a detailed analysis of its optical spectrum revealed the presence of a ‘standard’ accretion disk emission (as in FSRQs), whose light was probably swamped by the strong non-thermal continuum coming from the oriented relativistic jet (Padovani et al. 2019). This observation sets TXS 0506+056 apart from ‘true’ BL Lac objects that genuinely lack of an efficient accretion disk and a BLR, probably due to the low accretion rates of their central super massive black holes (SMBHs; e.g. see Hardcastle & Croston 2020), and suggests an emission mechanism that is similar to that considered for FSRQs (Keivani et al., 2018).
@@
The search for positional coincidences between neutrino events (or excesses of events) and extragalactic sources is hampered by the large positional uncertainties of the events. These uncertainties make the number of chance coincidences very high, even for relatively rare objects like blazars.
@@
Hints of positional correlations between neutrino events and BL Lac objects have been found by Padovani et al. (2016) and Giommi et al. (2020b). Other works have suggested a possible association between IceCube events and the most radio-luminous blazars (150 mJy), including both FSRQs and BL Lacs Plavin et al. 2020, 2021, 2023; Buson et al. 2022, 2023.
@@
Even if relativistic jets are natural candidates for neutrino emission, the recent association of a radio-quiet (RQ) AGN (NGC1068) with neutrino events (IceCube Collaboration et al. 2022) suggests that other mechanisms are at work. For instance, Abbasi et al. (2022a) have shown that the acceleration of cosmic rays responsible for the neutrino emission can occur in the accretion disk region of AGNs. The multi-wavelength analysis of NGC 1068 conducted within a p- scenario indicates that the most plausible location for the neutrino source is in close proximity to the SMBH, which is embedded in an extremely photon-dense environment (Murase, 2022; Padovani et al., 2024). This finding suggests the AGN corona as the likely origin, while star-forming regions, the jet, and the outflow are disfavoured. Yang et al. (2025) find that even in the case of TXS 0506+056, the accretion flow might explain both the long-term steady neutrino emission and the outburst, for which a super Eddington accretion is required. For this reason, the simple positional match of a sample of blazars with neutrino events, even when statistically significant, does not necessarily imply that the emission is due to the relativistic jets, since some of these sources (the FSRQs) also have a powerful accretion disk similar to that observed in RQ quasars.
@@
In this paper we test the hypothesis that FSRQs are responsible, at least in part, for the astrophysical neutrino events detected by IceCube. At the same time, we aim to establish whether this is really due to the presence of an oriented jet or, instead, can be associated with the RQ component of a FSRQ, i.e. the accretion disk and corona. To this end, we started from an optically selected sample of AGNs (that includes both radio-loud and RQ objects), and from this we extracted both the FSRQ and the RQ quasar samples. From an analysis of their correlation with neutrino events, we assessed not only the association of neutrino production with quasars, but also the possible origin site of the neutrinos (jets or disk-corona accretion).
@@
In Sect. 2 we present the quasar data sample and the list of IceCube events that we considered in the analysis. In Sect. 3 we describe the positional quasar–neutrino cross-match and present the results for both FSRQs and RQ quasars. In Sect. 4 we analyse the distribution of the minimum distance between neutrinos and quasars. In Sect. 5 the most likely neutrino emitter candidates are presented individually, and in Sect. 6, we discuss the purity of the sample together with the estimated fraction of high-energy neutrino events produced by FSRQs. Finally, in Sect. 7 we put the results in the context of recent findings presented in the literature.
@@
Throughout the paper we assume a flat cold dark matter cosmology with H0=71 km s-1 Mpc-1, =0.7, and =0.3. Spectral indices are given assuming S.
2 Quasar and neutrino catalogues
2.1 The quasar samples
@@
We started with the sample of type 1 AGNs presented in Shen et al. (2011, hereafter the S11 sample), which is based on the data from the Sloan Digital Sky Survey (SDSS). The sample includes 105,783 spectroscopically confirmed type 1 AGNs, mostly in the northern hemisphere, and contains several pieces of information on the central engine, such as the mass of the SMBH and its accretion rate, which can be used to further investigate potential neutrino emitters. For our analysis we only considered the sky area within 120 R.A. 260∘ and 0∘ Dec. 60∘, where the coverage is uniform (Fig. 1); this area includes a total of 89,437 quasars.
@@
For the radio information we used the Cosmic Lens All Sky Survey (CLASS; Myers et al. 2003; Browne et al. 2003), which homogeneously covers all the extragalactic northern sky. The CLASS catalogue contains 11,000 flat spectrum radio sources selected at 5 GHz down to a well-defined flux density limit (30 mJy). Since blazars are flat spectrum radio sources, this survey represents the best starting point for our selection. The survey is also deep enough to select most of the blazars with an optical magnitude brighter than 20 (which roughly corresponds to the limit of the S11 sample). Thanks to the excellent positional accuracy of the survey (1″) the optical/radio association can be performed with high reliability (Fig. 1). The cross-correlation between the S11 sample and CLASS produces 1218 sources (1122 in the region of interest), using 1″ positional tolerance. We call this sample CLASH, and throughout the rest of the paper, unless otherwise specified, flux densities are measured at 5 GHz.
@@
For comparison, we also built a sample of RQ quasars, defined as those sources from S11 not detected in CLASS and located more than 5 from any source in the FIRST catalogue. FIRST is a radio survey at 1.4 GHz with flux density limit of 1 mJy, significantly deeper than CLASS and with a good positional accuracy ( 2). Given the typical magnitudes of the Shen et al. (2011) quasars, the non-detection of a source in the FIRST survey implies a radio-loudness222The radio-loudness is defined as the ratio between the radio and the optical monochromatic luminosities (per unit of frequency) computed, respectively, at 5 GHz and at 4400Å (Kellermann et al. 1989) below 10 i.e. in the typical RQ regime (Kellermann et al. 1989). Therefore, the RQ quasar sample should mostly contain non-jetted objects. This control sample comprises 73,339 RQ objects.
@@
As shown in Fig. 2, the two populations are indistinguishable in terms of redshift, mass, bolometric luminosity and Eddington ratio. The only difference is the presence of a strong radio emission, which we assume is produced by a relativistic jet oriented towards our line of sight.
@@ 2.2 The IceCube neutrino events
For our study we used the IceCube Event Catalog of Alert Tracks (IceCat-1; Abbasi et al. 2023b), which includes 267 non-vetoed track-like neutrino events likely produced by astrophysical sources between 2011 and 2021. Since the positional uncertainty is extremely large in a significant fraction of the events, we considered only those events with the best positional accuracy:
@@
We stress that these requirements were frozen at the beginning of the analysis, and no optimization has been performed. In the region of interest defined above (Fig. 1), these constraints produce a list of 15 events (Table 1), with IC121103A, IC190503, and IC200117A excluded because too close to the border. Following the approach of Plavin et al. (2023), also discussed in Abbasi et al. (2023a), the 90% error region is assumed to be encircled by the 4 elliptical segments with the RA and Dec. errors as semi-axes. The energy distribution of the 15 events is shown in Fig. 3: the selection on the positional accuracy includes events belonging to the bulk of the population detected in the northern sky with energies in the 100-500 TeV interval (Table 1).
3 Quasar–neutrino associations
3.1 CLASH FSRQ–neutrino association
@@
Considering the list of 15 IceCat-1 events and the 1122 CLASH FSRQ we find that 9 CLASH sources fall within the error boxes of 7 IceCube events (Table 1). Varying the flux density limit of the CLASH sample, the number of objects varies from 1122 with a limit of 30 mJy to 270, when the flux density limit is 200 mJy. The number of neutrino events, with at least one CLASH quasar within the 90% error region varies from 7 to 3, while the number of CLASH included in the neutrino 90% confidence error region varies from 9 to 3 (Table 2).
@@
Multiple matches clearly indicate that the probability of chance coincidence is not negligible. We estimated the expected number of chance coincidences by randomly shifting quasar positions within a circular annulus of radius 3–10∘ around their original locations. This procedure was repeated 105 times for different radio flux density limits of the FSRQ sample. The statistical significance of the observed correlation (p-value) was calculated by counting the realizations in which the number of neutrinos including at least one CLASH FSRQ exceeded that observed in the actual dataset. The output of this Monte Carlo simulation at different radio flux density limits are reported in Fig. 4 and Table 2.
@@
These values show a minimum located at a flux density limit of 100 mJy (Fig. 4). At this flux density limit, the number of CLASH FSRQs is 497, with 7 of them in the error box of 6 neutrino events. The average number of associations of an equal number of random positions with the 15 neutrino error box is 1.3. We find that in 52 cases out of 1e5 an equal or larger number neutrino have a CLASH in the error box. Considering the number of different trials involved in this analysis (6 with different flux density limits), we estimate that the final probability corresponds to a significance of 2.7. We obtain identical results when repeating the simulation by randomly shifting the positions of the neutrinos instead of the quasars.
| fmin | Quasars | #n | #n | p-val | KS | KS (20∘) |
| 30 | 1122 | 7 | 2.6 | 6.3e-03 | 2.8e-02 | 8.2e-05 |
| 75 | 614 | 6 | 1.4 | 1.7e-03 | 6.5e-03 | 7.8e-06 |
| 100 | 497 | 6 | 1.2 | 5.3e-04 | 3.3e-03 | 4.8e-06 |
| 125 | 402 | 5 | 0.9 | 1.5e-03 | 2.4e-03 | 9.4e-06 |
| 175 | 300 | 3 | 0.7 | 3.0e-02 | 8.4e-03 | 5.3e-06 |
| 200 | 270 | 3 | 0.6 | 2.3e-02 | 2.9e-02 | 1.4e-04 |
3.2 RQ quasar–neutrino association